# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ D.5.2 Strongly convex case
$\mathbb{E}\left[\ell(\theta_{T};(v,-\frac{C}{\sqrt{L}}))-\ell(\theta^{\prime}_{T};(v,-\frac{C}{\sqrt{L}}))\right]=C\sqrt{L}\mathbb{E}\lambda_{T}+\frac{C^{2}}{2L}-\frac{C^{2}}{2L}\in\Omega\left(pC^{2}\gamma T\right)$
Appendix E Deferred proofs and additional details from Section 4
We provide details supporting Section 4, including numerical experiments E.1 and generalization error computations E.2.
E.1 Numerical analysis additional details
@@
We specify here the details of our numerical experiment presented in Section 4 into three points: (i) the problem instanciation; (ii) our tailored Byzantine failure attack; and (iii) how we compute the empirical robustness parameter.
@@
    We instantiate the setting from the proof of theorem D.13 (see full details therein). That is, we consider $\mathcal{A}=\mathrm{GD}$ under a linear loss function, using the parameter $C=1$, $\gamma=1$, $T=5$, $n=15$, $m=1$. We make vary $f$ from $1$ to $7$, and for each value instantiate the dataset presented in the proof of theorem D.13.
@@
    In our tailored Byzantine failure attack, the identities of the Byzantine workers remain fixed throughout the optimization process, but vary depending on the number of Byzantine workers $f$. The set of Byzantine identities is defined as :=$\mathcal{B}\vcentcolon=\{1,\ldots,n\}\setminus\mathcal{H}$, and is summarized in the table below. As we will see, we optimize the values sent by Byzantine workers so that they are selected while also biasing the gradient statistics—either toward positive or negative values—depending on the direction of the first parameter. To construct a maximally impactful scenario, we choose the identities of the Byzantine workers to maximize the minimal variance over all subsets of size $n-f$. This increases their biasing power, as it hinges on the smallest variance among these subsets. For instance, when $f=2$, the Byzantine workers can be hidden in the dominant subset $N$; however, when $f=6$, they must be distributed across both competing subgroups $E$ and $F$ to maintain influence.
@@
    The Byzantine strategy mimics the poisoning attack in the first iteration, then amplifies the resulting parameter direction by adaptively crafting updates that remain within the aggregation rule’s selection range. Specifically, we describe their algorithm in what follows:
    • If $\theta=0$, all Byzantine workers send the same value as the one in the poisoning attack. This value depends on each worker’s identity; see the proof of Theorem D.13 for details.
    • If $\theta>0$, all Byzantine workers send the value $-\alpha_{f}$, where $\alpha_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
    • If $\theta<0$, all Byzantine workers send the value $\beta_{f}$, where $\beta_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
@@
    if $\theta=0$ then
           return $g^{(i)}_{\text{poisoning}}$# the same value as data poisoning
    else if $\theta>0$ then
@@
  • (iii)
    Let denote $\hat{\kappa}_{f},\hat{\kappa}^{\prime}_{f}$ the empirical measures of the robustness coefficients observed over the optimization trajectories. The empirical robustness coefficients are computed as follows:
    $\hat{\kappa}_{f}=\max_{t\in\{0,\ldots,T-1\}}\max_{\begin{subarray}{c}S\subset\{1,\ldots,n\}\\ |S|=n-f\end{subarray}}\left(\frac{{|\overline{g}_{S_{t}^{*}}-\overline{g}_{S}|}^{2}}{\frac{1}{|S|}\sum_{i\in S}{\left(g^{(i)}_{t}-\overline{g}_{S}\right)}^{2}}\right),$
    and
    $\hat{\kappa}^{\prime}_{f}=\max_{t\in\{0,\ldots,T-1\}}\max_{\begin{subarray}{c}S\subset\{1,\ldots,n\}\\ |S|=n-f\end{subarray}}\left(\frac{{|\overline{g}^{\prime}_{S^{\prime*}_{t}}-\overline{g}^{\prime}_{S}|}^{2}}{\frac{1}{|S|}\sum_{i\in S}{\left(g^{\prime(i)}_{t}-\overline{g}^{\prime}_{S}\right)}^{2}}\right),$
    where $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g^{(i)}_{t}$ and $S_{t}^{*}$ the subset chosen by the $\mathrm{SMEA}$ aggregation rule at iteration $t$. The same notation applies to the other run. To assess the gap between our theoretical upper bound and its empirical realization, we analyze the tightness of the robustness coefficient $\hat{\kappa}_{f}$ as observed in our setup. We find that $\hat{\kappa}_{f}$ is consistently smaller than its theoretical worst-case value $\kappa_{\mathrm{SMEA}}=\frac{4f}{n-f}{\left(1+\frac{f}{n-2f}\right)}^{2}$. Substituting this empirical value into our theoretical upper bound, $\gamma CT\left(\frac{2}{(n-f)m}+2\sqrt{\kappa}\right)$, significantly reduces the discrepancy between theoretical predictions and actual performance:
    $\gamma CT\left(\frac{2}{(n-f)m}+\sqrt{\hat{\kappa}_{f}}+\sqrt{\hat{\kappa}^{\prime}_{f}}\right).$
E.2 Details on the computation of the generalization error
We again instantiate the problem outlined in the proof of our lower bound (Theorem 3.3, see Section D.5.1), under data poisoning attack, with $\gamma=C=1$ and $m=1$. To calculate the generalization error, we assume every honest local distribution is a Dirac with singularity at the value defined in the proof of Theorem 3.3—except for a “pivot” worker (indexed $1$ without loss of generality) whose distribution is a mixture $p_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{-1}$. In this case, the two possible datasets are always neighboring. We denote by $\theta_{T}^{(0)}$ and $\theta_{T}^{(-1)}$ the parameters output by the robust distributed $\mathrm{GD}$ algorithm with $\mathrm{SMEA}$ on the two different datasets. We denote by $\mathcal{S}$ a dataset sampled from the distribution specified above, and let $z_{1}$ be the sample held by worker 1. In this setup, we are able to compute the generalization error exactly as follows:
$=\frac{1}{|\mathcal{H}|}\mathbb{E}_{\mathcal{S}}\big{[}\frac{1}{2}\theta_{T}^{(z_{1})}\times 0+\frac{1}{2}\theta_{T}^{(z_{1})}\times(-1)-\theta_{T}^{(z_{1})}z_{1}\big{]}$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{1}{2}\mathbb{E}_{\mathcal{S}}\theta^{(z_{1})}_{T}-\mathbb{E}_{\mathcal{S}}\big{[}\theta_{T}^{(z_{1})}z_{1}\big{]}\right)$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{1}{2}\left(\frac{1}{2}\theta^{(0)}_{T}+\frac{1}{2}\theta^{(-1)}_{T}\right)-\frac{1}{2}\theta^{(0)}_{T}\times 0-\frac{1}{2}\theta^{(-1)}_{T}\times(-1)\right)$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{1}{4}\theta^{(0)}_{T}-\frac{1}{4}\theta^{(-1)}_{T}+\frac{1}{2}\theta^{(-1)}_{T}\right)$
$=\frac{\theta^{(-1)}_{T}-\theta^{(0)}_{T}}{4|\mathcal{H}|}$
Hence:
$\mathbb{E}_{\mathcal{S}}\big{[}R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big{]}=\frac{\theta^{(-1)}_{T}-\theta^{(0)}_{T}}{4(n-f)}.$
This value corresponds to the stability plots shown in Figure 1, modulo the factor $\frac{1}{4(n-f)}$. Hence, in the wort-case, the generalization error under Byzantine failure and data poisoning attacks exhibits the same discrepancy as the one observed in uniform algorithmic stability.
@@ Appendix F Refined bounds under bounded heterogeneity and bounded variance assumptions
As shown in the following, our analysis techniques can yield tighter bounds when assuming bounded heterogeneity and bounded variance. However, we note that these bounds do not provide additional conceptual insights within the scope of our discussion. Indeed, within the uniform stability framework, if we assume only bounded gradients and nothing more, there exist distributed learning settings where $G$ is a constant multiple of the Lipschitz constant $C$. For instance, consider a scenario where half of the honest workers produce gradients with norm $C$, and the other half produce zero gradients; in this case, $G$ would be $C/2$.
As an example, we derive bounds on the expected trace and spectral norm of the empirical covariance matrix of honest workers’ gradients that are expressed in terms of $G$ and $\sigma$. These estimates can be directly used in Theorems D.3 and D.5 to provide tighter bounds under additional assumptions of bounded heterogeneity and bounded variance.
Lemma F.1 (Expectation bound for the empirical covariance trace).
Consider the setting described in Section 2 under Byzantine attacks with Assumptions F.1 and F.2. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa,\operatorname{Tr})$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
$\mathbb{E}_{\mathcal{A}}\operatorname{Tr}\Sigma_{\mathcal{H},t}\leq 3G^{2}+3\sigma^{2}(1+\frac{1}{n-f})$
Proof.
We decompose the sum into the following:
$\mathbb{E}_{\mathcal{A}}\operatorname{Tr}\Sigma_{\mathcal{H},t}$ $=\mathbb{E}\bigg{[}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|g_{t}^{(i)}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}^{(i)}\|_{2}^{2}\bigg{]}$
$\leq\frac{3}{n-f}\sum_{i\in\mathcal{H}}\mathbb{E}[\|g_{t}^{(i)}-\nabla\widehat{R}_{i}(\theta_{t})\|_{2}^{2}]+3\mathbb{E}[\frac{1}{n-f}\sum_{i\in\mathcal{H}}\|\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\|_{2}^{2}]$
$\quad\quad+\frac{3}{n-f}\sum_{i\in\mathcal{H}}\mathbb{E}[\|\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}^{(i)}\|_{2}^{2}]$
$\leq 3G^{2}+3\sigma^{2}+\frac{3}{{(n-f)}^{2}}\sum_{i\in\mathcal{H}}\mathbb{E}[\|g_{t}^{(i)}-\nabla\widehat{R}_{i}(\theta_{t})\|_{2}^{2}]$
Lemma F.2 (Expectation bound for the empirical covariance spectral norm).
Consider the setting described in Section 2 under Byzantine attacks with Assumptions F.1 and F.2. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa,\|\cdot\|_{sp})$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
@@
and using Lemma D.4 from [5]:
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}2\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle\right]\\ \quad=\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\langle v,Mv\rangle\right]\hfill\\ \quad\leq 9^{d}\sup_{\|v\|_{2}\leq 1}\mathbb{E}\left[\langle v,Mv\rangle\right]\hfill\\ \quad=9^{d}\sup_{\|v\|_{2}\leq 1}2\mathbb{E}\sum_{i\in\mathcal{H}}\langle v,\underbrace{\mathbb{E}[g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}|\theta_{t}]}_{=0}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle$
@@
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle^{2}\right]\\ \leq\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\rangle^{2}\right]\\ \leq\mathbb{E}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\underbrace{\mathbb{E}_{J^{(i)}_{t}}[\|g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\|^{2}_{2}]}_{\leq\sigma^{2}}\right]$
@@
Appendix G High-probability excess risk bounds for robust distributed learning under smooth and strongly convex loss functions
In this section, we establish high-probability excess risk bounds (see Equation 1) under smooth and strongly convex assumptions for robust distributed learning algorithms, providing insights into their generalization performance. Specifically, we combine a deterministic optimization error bound—adapted directly from prior work [15]—with our high-probability stability bound.
High probability bound under uniform stability framework are derived as follows. The change of a single training sample has a small effect on a uniformly stable algorithm. This broadly implies low variance in the difference between empirical and population risk. If this difference also has low expectation, the empirical risk should accurately approximate the population risk [10]. In fact, we directly adapt the results from [31] to our framework and state the following high-probability bound.

### v2

@@ D.2.2 Strongly convex case
$\mathbb{E}\left[\ell(\theta_{T};(v,-\frac{C}{\sqrt{L}}))-\ell(\theta^{\prime}_{T};(v,-\frac{C}{\sqrt{L}}))\right]=C\sqrt{L}\mathbb{E}\lambda_{T}+\frac{C^{2}}{2L}-\frac{C^{2}}{2L}\in\Omega\left(pC^{2}\gamma T\right)\qed$
Appendix E DEFERED PROOFS AND ADDITIONAL DETAILS FROM SECTION 4
E.1 Proof of Lemmas 4.1 and 4.2
Below we compute the generalization error of a learning algorithm in the setting explicited in Section 4.
Lemma 4.1.
Consider the setting described in Section 2, with $m=1$, regardless of the assumed threat model. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$, such that for any distributed algorithm $\mathcal{A}$, we have:
$\left|\mathbb{E}_{\mathcal{A},\mathcal{S}\sim\otimes_{i\in\mathcal{H}}\left(p_{i}^{\otimes m}\right)}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]\right|=\frac{1}{4(n-f)}\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}\left[\ell(\mathcal{A}(S);z)-\ell(\mathcal{A}(S^{\prime});z)\right],$ (16)
where the dependence on the threat model arises solely through the stability term.
Proof.
We consider a linear loss function $z\in[-C,C],\theta\in\mathbb{R}\mapsto z\theta$, and we assume every honest local distribution is a Dirac (regardless of the value where the singularity is), except for a “pivot” honest worker (indexed $1$ without loss of generality) whose distribution is a mixture of Dirac $p_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{-C}$. In this case, when the number of local samples is fixed to one ($m=1$), the two possible datasets are always neighboring. Denote them $S^{(-C)}$ and $S^{(0)}$. We denote the parameters resulting from $\mathcal{A}$ on the two different datasets by $\mathcal{A}(S^{(0)})=\theta_{T}^{(0)}$ and $\mathcal{A}(S^{(-C)})=\theta_{T}^{(-C)}$.
In this setup, we can compute the generalization error exactly as follows:
$=\frac{1}{|\mathcal{H}|}\mathbb{E}_{\mathcal{S}}\big[\frac{1}{2}\theta_{T}^{(z_{1})}\times 0+\frac{1}{2}\theta_{T}^{(z_{1})}\times(-C)-\theta_{T}^{(z_{1})}z_{1}\big]$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{C}{2}\mathbb{E}_{\mathcal{S}}\theta^{(z_{1})}_{T}-\mathbb{E}_{\mathcal{S}}\big[\theta_{T}^{(z_{1})}z_{1}\big]\right)$
$=\frac{1}{|\mathcal{H}|}\left(-\frac{C}{2}\left(\frac{1}{2}\theta^{(0)}_{T}+\frac{1}{2}\theta^{(-C)}_{T}\right)-\frac{1}{2}\theta^{(0)}_{T}\times 0-\frac{1}{2}\theta^{(-C)}_{T}\times(-C)\right)$
$=\frac{C}{|\mathcal{H}|}\left(-\frac{1}{4}\theta^{(0)}_{T}-\frac{1}{4}\theta^{(-C)}_{T}+\frac{1}{2}\theta^{(-C)}_{T}\right)$
$=\frac{C\left(\theta^{(-C)}_{T}-\theta^{(0)}_{T}\right)}{4(n-f)}$ (17)
As the above computation is independent of the assumed threat model, we have,
$\left|\mathbb{E}_{\mathcal{A},\mathcal{S}}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]\right|=\frac{1}{4(n-f)}\sup_{z\in[-C,C]}\mathbb{E}_{\mathcal{A}}\left|\ell(\mathcal{A}(S^{(-C)});z)-\ell(\mathcal{A}(S^{(0)});z)\right|.\qed$
Building upon the theorem above, we leverage our stability analysis to demonstrate that the Byzantine failure and data poisoning threat models differ fundamentally in their capacity to affect the generalization error of distributed learning algorithms.
Theorem 4.2.
Consider the setting of Lemma 4.1, $\frac{n}{3}\leq f<\frac{n}{2}$, $\mathcal{A}=\mathrm{GD}$ with $\mathrm{SMEA}$. There exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, data distributions ${\{p_{i}\}}_{i\in\mathcal{H}}$ over $\mathcal{Z}$ and a Byzantine attack such that for any data poisoning attacks, we have
$\frac{\mathcal{E}^{\mathrm{byz}}_{gen}}{\mathcal{E}^{\mathrm{pois}}_{gen}}\in\Omega\left(\frac{n-f}{\sqrt{f(n-2f)}}\right),$ (18)
where $\mathcal{E}^{\mathrm{byz}}_{gen}$ and $\mathcal{E}^{\mathrm{pois}}_{gen}$ denote generalization errors under the Byzantine and data poisoning attacks, respectively.
Proof.
We specialize Lemma 4.1 to $\mathcal{A}=\mathrm{GD}$ with $\mathrm{SMEA}$. If we assume every honest local distribution is a Dirac with singularity at the value defined in the paragraph (i) of the proof of Theorem D.2—except for a “pivot” honest worker, indexed $1$, whose distribution is a mixture $p_{1}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{-C}$, then the only two possible datasets for honest workers align with the worst-case Byzantine and data-poisoning scenarios considered in our lower bounds (see Sections D.2.1 and C.3 for details of the datasets used). Hence, using the attacks described therein demonstrates that the generalization error exhibits the same intrinsic gap between the two threat models, consistent with the uniform stability analysis, as illustrated in Figure 1. That is $\mathcal{E}^{\mathrm{byz}}_{gen}$ scales according to eq. 4 and $\mathcal{E}^{\mathrm{pois}}_{gen}$ scales according to eq. 7 (both scaled by $\frac{1}{4\left(n-f\right)}$), as revealed by section E.1. ∎
E.2 Numerical analysis additional details
@@
We describe here the details of our numerical experiment presented in Section 4 into two points: (i) the problem instantiation; (ii) our tailored Byzantine attack.
@@
    We instantiate the setting from the proof of Theorem D.2 (see full details therein). That is, we consider $\mathcal{A}=\mathrm{GD}$ under a linear loss function, using the parameters $C=1$, $\gamma=1$, $T=5$, $n=15$, $m=1$. We make vary $f$ from $1$ to $7$, and for each value instantiate the dataset presented in the proof of Theorem D.2.
@@
    In our tailored Byzantine failures, the identities of the Byzantine workers remain fixed throughout the optimization process, but vary depending on the number of Byzantine workers $f$. The set of Byzantine identities is defined as $\mathcal{B}\vcentcolon=\{1,\ldots,n\}\setminus\mathcal{H}$, and is summarized in the table below. As we will see, we optimize the values sent by Byzantine workers so that they are selected while also biasing the gradient statistics—either toward positive or negative values—depending on an event that triggers them to switch their communicated value. To construct a maximally impactful scenario, we choose the identities of the Byzantine workers to maximize the minimal variance over all subsets of size nfn-f. This increases their biasing power, as it hinges on the smallest variance among these subsets. For instance, when $f=2$, the Byzantine workers can be hidden in the dominant subset $N$; however, when $f=6$, they must be distributed across both competing subgroups $E$ and $F$ to maintain influence.
@@
    For instance, if a single sample from worker $1$ (without loss of generality) is changed from 0 to C-C (defining a neighboring dataset), the server receives a different average $\frac{1}{m}C<0$ that triggers the Byzantine workers to amplify the positive or negative parameter direction by adaptively crafting updates that remain within the aggregation rule’s selection range. Specifically, their algorithm is as follows:
    • If $g_{0}^{(1)}\geq 0$, all Byzantine workers send the value $\alpha_{f}$, where $\alpha_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
    • If $g_{0}^{(1)}<0$, all Byzantine workers send the value $\beta_{f}$, where $\beta_{f}$ is numerically chosen to be as large as possible under the constraint that the $\mathrm{SMEA}$ aggregation rule continues to select the values sent by Byzantine workers in subsequent steps.
@@
    if $g_{0}^{(1)}\geq 0$ then
@@ Appendix F REFINED BOUNDS UNDER BOUNDED HETEROGENEITY AND BOUNDED VARIANCE ASSUMPTIONS
As shown in the following, our analysis techniques can yield tighter bounds when assuming bounded heterogeneity and bounded variance. However, we note that these bounds do not provide additional conceptual insights within the scope of our discussion. Indeed, within the uniform stability framework, if we assume only bounded gradients and nothing more, there exist distributed learning settings where $G$ is a constant multiple of the Lipschitz constant $C$. For instance, consider a scenario where half of the honest workers produce gradients with norm $C$, and the other half produce gradients with norm C-C; in this case, $G$ would be $C$.
As an example, we derive bounds on the expected spectral norm of the empirical covariance matrix of honest workers’ gradients that are expressed in terms of $G$ and $\sigma$. These estimates can be directly used in Theorem 3.1 to provide tighter bounds under additional assumptions of bounded heterogeneity and bounded variance.
Lemma F.1.
Consider the setting described in Section 2 under Byzantine attacks with Assumptions F.1 and F.2. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
@@
and using Lemma D.4 from Allouah et al. (2023b):
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}2\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle\right]\\ \quad=\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\langle v,Mv\rangle\right]\leq 9^{d}\sup_{\|v\|_{2}\leq 1}\mathbb{E}\left[\langle v,Mv\rangle\right]\hfill\\ \quad=9^{d}\sup_{\|v\|_{2}\leq 1}2\mathbb{E}\sum_{i\in\mathcal{H}}\langle v,\underbrace{\mathbb{E}[g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}|\theta_{t}]}_{=0}\rangle\langle v,\nabla\widehat{R}_{i}(\theta_{t})-\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})\rangle$
@@
$\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})+\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})-\overline{g}_{t}\rangle^{2}\right]$ $\leq\mathbb{E}\left[\sup_{\|v\|_{2}\leq 1}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\langle v,g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\rangle^{2}\right]$
$\leq\mathbb{E}\left[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\underbrace{\mathbb{E}_{J^{(i)}_{t}}[\|g^{(i)}_{t}-\nabla\widehat{R}_{i}(\theta_{t})\|^{2}_{2}]}_{\leq\sigma^{2}}\right]$
@@
Appendix G HIGH-PROBABILITY EXCESS RISK BOUNDS FOR ROBUST DISTRIBUTED LEARNING
In this section, we establish high-probability excess risk bounds (see Equation 1) under smooth and strongly convex assumptions for robust distributed learning algorithms, providing insights into their generalization performance. Specifically, we combine a deterministic optimization error bound—adapted directly from prior work from Farhadkhani et al. (2024b)—with our high-probability stability bound.
High probability bound under uniform stability framework are derived as follows. The change of a single training sample has a small effect on a uniformly stable algorithm. This broadly implies low variance in the difference between empirical and population risk. If this difference also has low expectation, the empirical risk should accurately approximate the population risk (Bousquet and Elisseeff, 2002). In fact, we directly adapt the results from Bousquet et al. (2020) to our framework and state the following high-probability bound.